Attempting to assist my son with homework. topic is "order of operations". Here's the math problem......2+7^2-3^2/3^2-1x5=35. Here's how we did it.... 2+49-9/9-1x5=35
51-9/9-1x5=35
51-1-1x5=35
50-1x5=35
49x5=35????????

Answer by Flannery(124)

Remember to "Please Excuse My Dear Aunt Sally".
Do PARENTHESES first from innermost to outer most and then from left to right.
Then do EXPONANTS left to right.
Then MULTIPLY left to right.
Then DIVIDE left to right.
Then ADD left to right.
Finally SUBTRACT left to right.
2+7^2-3^2/3^2-1*5=
2+49-9/9-1*5=
2+49-9/9-5= (-1*5=-5)
2+49-1-5= (-9/9=-1)
51-1-5=
50-5=
45
I checked this solution on Microsoft Mathematics.